01) Identify the X intercepts of this quadratic function y = 2x- x²


(02) Find the coordinates of maximum point of y = -4x + 8x – 2 by using equation y by using equetion method.



(03) Graph the quadratic function y = -3x +x+1 and you are required to answer the followings by using the graph:



(i) Coordinates of maximum point


(ii) Axis of symmetry


1
Expert's answer
2022-01-11T01:04:46-0500

(01)

X-intercept(s)


"y=0=> 2x- x^2=0"

"x(2-x)=0"

"x_1=0, x_2=2"

X-intercept(s): "(0, 0), (2, 0)."


(02)


"y=-3x^2+8x-2"

"y=-3(x^2-2(\dfrac{4}{3})x+(\dfrac{4}{3})^2)+\dfrac{16}{3}-2"

"y=-3(x-\dfrac{4}{3})^2+\dfrac{10}{3}"

The maximum point is "(\dfrac{4}{3}, \dfrac{10}{3})."


(03)


"y = -3x^2 +x+1""y=-3(x^2-2(\dfrac{1}{6})x+(\dfrac{1}{6})^2)+\dfrac{1}{12}+1""y=-3(x-\dfrac{1}{6})^2+\dfrac{13}{12}"

The maximum point is "(\dfrac{1}{6}, \dfrac{13}{12})."


Axis of symmetry: "x=\dfrac{1}{6}"

y-intercept


"x=0=> y=-3(0)^2+0+1=1"

y-intercept: "(0, 1)."

X-intercept(s)


"y=0=> -3x^2+x+1=0""D=(1)^2-4(-3)(1)=13""x_{1,2}=\dfrac{-1\pm \sqrt{13}}{2(-3)}"

"x_{1}=\dfrac{1- \sqrt{13}}{6}, x_{2}=\dfrac{1+\sqrt{13}}{6}"

X-intercept(s): "(\dfrac{1- \sqrt{13}}{6}, 0), (\dfrac{1+\sqrt{13}}{6}, 0)."



(i) The maximum point is "(\dfrac{4}{3}, \dfrac{10}{3})."


(ii) Axis of symmetry: "x=\dfrac{1}{6}"